# JWST detection of a carbon dioxide dominated gas coma surrounding interstellar object 3I/ATLAS

### v1

@@ Abstract
3I/ATLAS is the third confirmed interstellar object to visit our Solar System, and only the second to display a clear coma. Infrared spectroscopy with the James Webb Space Telescope (JWST) provides the opportunity to measure its coma composition and determine the primary activity drivers. We report the first results from our JWST NIRSpec campaign for 3I/ATLAS, at an inbound heliocentric distance of au. The spectral images (spanning – m) reveal a CO2 dominated coma, with enhanced outgassing in the sunward direction, and the presence of H2O, CO, OCS, water ice and dust. The coma CO2/H2O mixing ratio of is among the highest ever observed in a comet, and is above the trend as a function of for long-period and Jupiter-family comets (excluding the outlier C/2016 R2). Our observations are compatible with an intrinsically CO2-rich nucleus, which may indicate that 3I/ATLAS contains ices exposed to higher levels of radiation than Solar System comets, or that it formed close to the CO2 ice line in its parent protoplanetary disk. A low coma H2O gas abundance may also be implied, for example, due to inhibited heat penetration into the nucleus, which could suppress the H2O sublimation rate relative to CO2 and CO.
@@ 1 Introduction
Comets and planetesimals are theorized to form in large numbers during the accretion of planetary systems. Many of these small bodies, composed of ice, rock, and dust, are subsequently expected to be ejected into interstellar space through gravitational encounters with larger (planetary or stellar) bodies (L. Dones et al., 2004; H. F. Levison et al., 2010; S. N. Raymond et al., 2018; S. Pfalzner & K. Vincke, 2020; X.-L. Zheng & J.-L. Zhou, 2025). The apparitions of the first confirmed interstellar objects (ISOs) 1I/‘Oumuamua in October 2017 and 2I/Borisov in August 2019 provided confirmation of this theory, and offered an unprecedented opportunity to study the nature of matter delivered to our Solar System from a distant planetary system, thus spawning a new field of planetary science.
Due to its faintness and short observing window, spectroscopic characterization of volatiles from 1I/‘Oumuamua proved elusive, so the composition of this object remains highly uncertain (‘Oumuamua ISSI Team et al., 2019). On the other hand, for the intrinsically brighter and more active 2I/Borisov, ultraviolet, optical, and submillimeter observations (e.g. M. A. Cordiner et al., 2020; D. Bodewits et al., 2020; P. Guzik & M. Drahus, 2021; C. Opitom et al., 2021; Z. Xing et al., 2020) provided tantalizing glimpses of its coma composition, and revealed an object that was similar in many ways to the well-studied comets from our own Solar System, but with an unusually strong enrichment in carbon monoxide (CO) (M. A. Cordiner et al., 2020; D. Bodewits et al., 2020). Considering the difficulty of studying the ices in the midplanes of protoplanetary disks and planetary systems elsewhere in our Galaxy, continued spectroscopic observations of interstellar objects have the potential to reveal crucial details on the physics and chemistry of planet formation in planetary systems other than our own.
@@
Similar to 2I/Borisov, 3I/ATLAS has been shown to display clear cometary activity (D. Jewitt & J. Luu, 2025; M. R. Alarcon et al., 2025; M. Minev et al., 2025). Early spectroscopic and photometric observations have revealed a compact nucleus (effective radius km), a bright, dusty coma, with a dust mass-loss rate of 12–120 kg s-1 (D. Jewitt et al., 2025a), and a red spectral slope, with possible water ice absorption at 2.0 µm (T. Kareta et al., in press; B. Yang et al., 2025). The Neil Gehrels-Swift Observatory detected ultraviolet emission from the gas-phase OH radical (Z. Xing et al., 2025), and assuming H2O to be the photolysis parent, a water production rate of s-1 was derived at a heliocentric distance of au.
Here, we present the first infrared spectroscopic observations from our campaign to observe 3I/ATLAS using the James Webb Space Telescope (JWST). This article focuses on analyzing the rovibrational fluorescence emission from H2O, CO2, and CO, which are the main drivers of coma activity in typical Solar System comets. Molecular production rates and mixing ratios are derived, enabling the first chemical characterization of the object’s volatile gas inventory.
@@ 2 Observations
Observations of 3I/ATLAS were performed using JWST (J. P. Gardner et al., 2023) on UT 2025-08-06 between 11:02–11:20, using the NIRSpec integral field unit (IFU), as part of program ID 5094. The IFU (T. Böker et al., 2022) divides its field of view into 30 parallel slices, each measuring . The light from each slice is then passed through a dispersing element (in this case, a fluorite prism), to produce 30 contiguous spectra on the NIRSpec detector. Subsequent image processing divides each IFU slice into 30 of spaxels, resulting in a spectral cube, covering the entire field of view. For reference, the FWHM of the JWST point-spread function is at 3 m. Each spectrum covers –5.3 m, with a resolving power that varies from at 1.2 m to at 5.3 m.
3I/ATLAS was acquired in the IFU and tracked using JPL Horizons ephemeris solution , which was constructed based on astrometry supplied by a large number of optical observers. The total on-target exposure duration was 640 s, divided across four dither positions, each spatially separated (in the approximate shape of a square) with offsets of from the (central) target position. The data were reduced using the JWST Calibration Pipeline software version v1.19.1 (H. Bushouse et al., 2025) using the JWST Calibration Reference Data System context file 1413. The four dithers were shifted and combined in the rest frame of the comet during image processing, thus allowing detector artifacts and cosmic rays to be identified and removed. A similar set of four exposures of the sky background was obtained, offset by along the horizontal axis of the IFU aperture. This allows contamination from background infrared sources, zodiacal light, and the telescope to be identified and subtracted. Data cubes with and without background subtraction were produced. The latter allows for the analysis of faint gas emission bands without the contribution of additional noise from a background subtraction. For each spectral data cube, the pipeline produces uncertainty and data quality maps, which were used during our analysis, in particular, for the derivation of formal () error estimates. After combining the four dithered observations, the absolute calibration accuracy is expected to be 3%.111https://jwst-docs.stsci.edu/jwst-calibration-status/nirspec-calibration-status/nirspec-ifu-calibration-status
@@ 3 Results
To isolate the CO2, H2O, and CO gas emission features, spectral data were extracted for each IFU pixel within the vicinity of each emission band, and a polynomial continuum fit was performed, excluding the spectral region directly inside each molecular emission band (see Figure 4). For CO2 and CO, a 3rd-order polynomial was used, whereas for H2O a 5th-order polynomial was used, to better fit the wing of the 3.0 m ice band (see Section 4). After continuum subtraction, the IFU spectra were integrated across the detected full emission width of each feature, then plotted as maps in Figure 2. The spatially averaged, continuum-subtracted emission band profile for each molecule is shown in the upper-right inset of their respective panels. The m scattered light image (integrated between 0.8–1.4 m) is also shown. The continuum brightness at 0.75 and 1.25 µm is 1013 and 1244 Jy inside a -diameter circular aperture, corresponding to cometary quantities (M. F. A’Hearn et al., 1984) of 392 and 492 cm, respectively (quoted without a correction for phase darkening).
@@
The gas and dust maps for 3I/ATLAS show a well-defined peak, offset East from the center of the IFU by . As shown by a later ephemeris reconstruction, about of the offset can be explained by a error on the predicted ephemeris, while the remaining offset remains under investigation. The 1.2 m scattered light and CO2 emission maps reveal an extended coma of dust and gas that spans the full extent of the NIRSpec IFU. The H2O and CO emission is weaker, and therefore noisier, but nevertheless confirms the presence of a spatially extended molecular coma. While the scattered light shows a clear asymmetry along the Sun-comet axis — enhanced in the direction of the sky-projected comet-Sun and velocity vectors (see also D. Jewitt et al. 2025b; C. O. Chandler et al. 2025) — the gas distributions appear relatively more symmetrical.
@@
To further investigate the coma structure, a enhanced version of the dust map (where is the sky-projected nucleocentric distance), is shown in Figure 3, and similarly for the gas maps in Appendix A (Figure 7). In these enhanced maps, the dominant coma spatial feature: dilution of the observed column densities due to quasi-spherical expansion, has been divided out. The -enhanced 1.2 m map reveals a strong, plume-like feature emanating from the pseudo-nucleus in the approximate direction of the Sun (slightly north of west), with an additional, weaker feature to the north-east. Since the dust in relatively faint comets such as 3I/ATLAS is optically thin at this wavelength, the shape of this feature is interpreted as resulting from an enhanced coma dust density in the sunward direction, potentially from the fragmentation of dust grains increasing the scattering cross section with distance from the nucleus. The gas maps on the other hand, show more subtle asymmetries. CO2 exhibits an azimuthal minimum towards the north, that we attribute to a combination of weak coma sub-structure and optical depth effects, since CO2 becomes optically thick close to the nucleus.
@@
Taking the ratio of continuum-subtracted fluxes within a circular aperture placed immediately adjacent to the sunward () and antisunward () sides of the brightest pixel (the assumed location of the nucleus) gives for the (1.2 m) dust, 1.07 for CO2, 1.04 for CO, and 1.20 for H2O (with uncertainties of on all measurements). Our observations thus reveal a heterogeneous coma morphology consistent with different outgassing patterns for the different species. Such heterogeneity can be explained by various factors, such as the different molecular sublimation temperatures, nucleus release and coma acceleration mechanisms, as well as differing inertial/fluid-dynamical properties for the gas and dust. The degree of asymmetry of the observed gases is likely related to their sublimation temperatures (), with , congruent with (M. Womack et al., 2017). This implies that the sublimation of CO2 and CO is more fully activated than H2O. Full interpretation of the observed outgassing morphologies, including the surprisingly strong sunward dust enhancement, will require detailed physical modeling. Nevertheless, our data are consistent with the origin of this dust feature being influenced by enhanced gas sublimation (and therefore, outgassing) rates associated with the higher temperature on the dayside of the nucleus.
@@
The 3.0 µm H2O ice band depth is 78.30.2% at 2.9–3.1 µm, with respect to the continuum at 2.5 and 3.8 µm. The band remains strong across the field of view, varying by no more than 10%. Water ice also has broad near-infrared absorption features at 1.5, 2.0, 3.0, and 4.5 µm. The 4.5 µm band is difficult to measure in our data due to blending with CO2 and CO gas emission and likely thermal continuum, the 2.0 µm band appears very weak, and there is no discernible 1.5 µm band. Furthermore, the 2.0 µm spectral region in our combined IFU dataset is affected by several bad pixels. In two dither positions with clean spectra in this region, the 2.0 µm band depth is (2.00.1)% and (1.60.1)%. This was estimated by normalizing the reflectance spectrum within a 04 radius aperture with a linear fit between 1.75 and 2.22 µm, and measuring the mean value at 1.95–2.05 µm. That the two dither positions are different by 3 suggests there are unaccounted-for systematic uncertainties in these estimates. Given the high-quality of the data and the water ice models of S. Protopapa et al. (2014), the appearance of a strong 3-µm band, but weak or absent 1.5- and 2.0-µm bands suggests micrometer or smaller ice grains. The apparent absence of the crystalline water ice Fresnel peak at 3.1 µm could suggest the presence of amorphous water ice, but the shape and strength of this feature depends on temperature (R. M. Mastrapa et al., 2009), and likely grain size and interaction with dust. Models of 120 K micrometer-sized amorphous and crystalline ice grains show little difference around the 3.1 µm region (S. Protopapa et al., 2014). Radiative transfer modeling is needed to further test the nature of the water ice grains in 3I/ATLAS.
@@ 4 Spectral Modeling
To derive gas production rates () and rotational temperatures (), the background-subtracted IFU data for CO2, CO, and H2O were subject to spectral modeling using optimal estimation routines as part of the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018, including synthetic fluorescence models described by G. Villanueva et al. 2025). Initially, we constructed spectral models by taking the average spectrum for each gas inside a circular aperture centered on the nucleus. The relatively high signal-to-noise ratio (SNR) in this IFU region assisted in helping define the choice of continuum shape and the gases to be included in the model. Figure 4 shows the integrated spectra within this aperture, along with the best-fitting PSG model gas emission components. For the m region, we also identified an emission band at 4.85 m consistent with OCS, which was included in the fit.
A zeroth-order polynomial fit was sufficient for describing the continua underneath the CO2 and CO features, whereas for H2O, we experimented with various analytic functions to produce a good fit, including the wing of the 3 m ice band. Although a 5th order polynomial was found to be sufficient to reproduce the shape of the continuum in this region, a lack of formal constraints across the 2.6–2.8 m region spanning the H2O 2.7 m band led to increased uncertainties on . Therefore, for H2O, we adopted a more physically constrained continuum model, formed as the product of a linear slope and a sigmoid function:
where is the linear continuum normalization point, is the linear slope, is the y-intercept point, is the depth of the 3-µm band, controls the slope of the band edge, and controls the wavelength of the band edge. The (variable) exponential onset of the sigmoid function effectively matches the quasi-Gaussian shape of the blue wing of the 3 m amorphous H2O ice band (e.g. G. Leto & G. A. Baratta, 2003).
The CO2 and CO spectral regions were modeled using the methods described by C. E. Woodward et al. (2025). The gas outflow velocity was set at 0.44 km s-1, based on the standard relationship (e.g. T. Ootsubo et al., 2012). Pixels close to the nucleus can be affected by significant line opacity and PSF-related flux losses, which are difficult to accurately model. Therefore, we subsequently proceeded with a “-curve” analysis, deriving the production rates and rotational temperatures as a function of distance from the nucleus within successive (independent) partial annular sectors centered on the brightest (nucleus-containing) pixel. To avoid flux losses from pixels at the very edge of the IFU, and to focus on a uniform coma angular region, a inscribed angle was used for all partial annuli, with radial bounds parallel to the north-east and south-east edges of the IFU; see inset diagram in Figure 5. We found a annulus width to provide sufficient signal-to-noise for all species. Given the very high SNR for CO2, we also generated a higher-resolution curve for this species using a annulus width. The resulting production rates and rotational temperatures as a function of are shown in Figure 5.
For the m region, H2O spectral models were generated using PSG, while the continuum was determined using Equation 1, combined with the =5000 Solar spectral model of R. C. Bohlin et al. (2014). Additional (weak) spectral contributions from coma CO2 to this region were added based on our best fit to the 4.3 m band. The H2O production rate and rotational temperature were retrieved using a Gaussian processes approach within the George software package (S. Ambikasaran et al., 2015), and uncertainties were derived using a Markov chain Monte Carlo algorithm (emcee; D. Foreman-Mackey et al., 2013).
@@
The curves for CO2 and CO (Figure 5) level off towards larger nucleocentric distances, indicating that gas production for these species is confined within km of the nucleus. On the other hand, the H2O curve shows no clear asymptote, and the error bars allow for the possibility of continued H2O production towards the edge of the NIRSpec IFU. Therefore, we cannot rule out a contribution to H2O gas in the outer coma from sublimating icy grains, which may be expected based on our detection of coma H2O ice. A strong icy-grain source of H2O between –10,000 km from the nucleus, would be required to reconcile our terminal value with the ultraviolet OH imaging measurements of Z. Xing et al. (2025), which indicated s-1 within a 20,000 km diameter aperture.
Although 13CO2 is securely detected in the nucleus-centered extract, optical depth effects preclude the derivation of a reliable 12CO2 production rate in this region. We attempted to retrieve (13CO2) in the annular sector between – from the nucleus, but only an upper limit s-1 could be obtained, resulting in a () lower limit on the C/13C ratio of , which is formally consistent with the terrestrial value of 89 (T. B. Coplen et al., 2002). See Appendix C for further details.

### v3

@@ Abstract
3I/ATLAS is the third confirmed interstellar object to visit our Solar System, and only the second to display a clear coma. Infrared spectroscopy with the James Webb Space Telescope (JWST) provides the opportunity to measure its coma composition and determine the primary activity drivers. We report the first results from our JWST NIRSpec campaign for 3I/ATLAS, at an inbound heliocentric distance of au. The spectral images (spanning – m) reveal a CO2 dominated coma, with enhanced outgassing in the sunward direction, and the presence of H2O, CO, water ice, dust and a tentative detection of OCS. The coma CO2/H2O mixing ratio of is among the highest ever observed in a comet, and is above the trend as a function of for long-period and Jupiter-family comets (excluding the outlier C/2016 R2). Our observations are compatible with an intrinsically CO2-rich nucleus, which may indicate that 3I/ATLAS contains ices exposed to higher levels of radiation than Solar System comets, or that it formed close to the CO2 ice line in its parent protoplanetary disk. A low coma H2O gas abundance may also be implied, for example, due to inhibited heat penetration into the nucleus, which could suppress the H2O sublimation rate relative to CO2 and CO.
@@ 1 Introduction
Comets and planetesimals are theorized to form in large numbers during the accretion of planetary systems. Many of these small bodies, composed of ice, rock, and dust, are subsequently expected to be ejected into interstellar space through gravitational encounters with larger, planetary or stellar, bodies (L. Dones et al., 2004; H. F. Levison et al., 2010; S. N. Raymond et al., 2018; S. Pfalzner & K. Vincke, 2020; X.-L. Zheng & J.-L. Zhou, 2025). The apparitions of the first confirmed interstellar objects (ISOs) 1I/‘Oumuamua in October 2017 and 2I/Borisov in August 2019 provided confirmation of this theory, and offered an unprecedented opportunity to study the nature of matter delivered to our Solar System from a distant planetary system, thus spawning a new field of planetary science.
Due to its faintness and short observing window, spectroscopic characterization of volatiles from 1I/‘Oumuamua proved elusive, so the composition of this object remains highly uncertain (‘Oumuamua ISSI Team et al., 2019). On the other hand, for the intrinsically brighter and more active 2I/Borisov, ultraviolet, optical, and submillimeter observations (e.g. M. A. Cordiner et al., 2020; D. Bodewits et al., 2020; P. Guzik & M. Drahus, 2021; C. Opitom et al., 2021; Z. Xing et al., 2020) provided intriguing glimpses of its coma composition, and revealed an object that was similar in many ways to the well-studied comets from our own Solar System, but with an unusually strong enrichment in carbon monoxide (CO) (M. A. Cordiner et al., 2020; D. Bodewits et al., 2020). Considering the difficulty of studying the ices in the midplanes of protoplanetary disks and planetary systems elsewhere in our Galaxy, continued spectroscopic observations of interstellar objects have the potential to reveal crucial details on the physics and chemistry of planet formation in planetary systems other than our own.
@@
Similar to 2I/Borisov, 3I/ATLAS has been shown to display clear cometary activity (D. Jewitt & J. Luu, 2025; M. R. Alarcon et al., 2025; M. Minev et al., 2025). Early spectroscopic and photometric observations revealed a compact nucleus (effective radius km), a bright, dusty coma, with a dust mass-loss rate of 12–120 kg s-1 (D. Jewitt et al., 2025), and a red spectral slope, with possible water ice absorption at 2.0 µm (T. Kareta et al., in press; B. Yang et al., 2025). The Neil Gehrels-Swift Observatory detected ultraviolet emission from the gas-phase OH radical (Z. Xing et al., 2025), and assuming H2O to be the photolysis parent, a water production rate of s-1 was derived at a heliocentric distance of au.
Here, we present the first set of infrared spectroscopic observations from our campaign to observe 3I/ATLAS using the James Webb Space Telescope (JWST). This article focuses on analyzing the rovibrational fluorescence emission from H2O, CO2, and CO, which are the main drivers of coma activity in typical Solar System comets. Molecular production rates and mixing ratios are derived, enabling chemical characterization of the object’s volatile gas inventory.
@@ 2 Observations
Observations of 3I/ATLAS were performed using JWST (J. P. Gardner et al., 2023) on UT 2025-08-06 between 11:02–11:20, using the NIRSpec integral field unit (IFU; T. Böker et al. 2022), as part of program ID 5094. The PRISM dispersive element was used, resulting in a array of spectra covering –5.3 m, with a resolving power that varies from at 1.2 m to at 5.3 m. The IFU pixel size is , which is approximately the same as the FWHM of the JWST point-spread function at 3 m.
3I/ATLAS was acquired and tracked in the IFU using JPL Horizons ephemeris solution , when the object was 2.73 au from the telescope, at au and a phase angle of . The total on-target exposure duration was 640 s, divided across four dither positions, each spatially separated (in the approximate shape of a square) with offsets of from the (central) targeted position. The data were reduced using the JWST Calibration Pipeline software version v1.19.1 (H. Bushouse et al., 2025) using the JWST Calibration Reference Data System context file 1413. The four dithers were shifted and combined in the rest frame of the comet during image processing, thus allowing detector artifacts and cosmic rays to be identified and removed. A similar set of four exposures of the sky background was obtained, offset by along the horizontal axis of the IFU aperture. This allows contamination from background infrared sources, zodiacal light, and the telescope to be identified and subtracted. Data cubes with and without background subtraction were produced. The latter allows for the analysis of faint gas emission bands without the contribution of additional noise from a background subtraction. For each spectral data cube, the pipeline produces uncertainty and data quality maps, which were used during our analysis, in particular, for the derivation of formal () error estimates. After combining the four dithered observations, the absolute calibration accuracy is expected to be 3%.111https://jwst-docs.stsci.edu/jwst-calibration-status/nirspec-calibration-status/nirspec-ifu-calibration-status
@@ 3 Results
To isolate the CO2, H2O, and CO gas emission features, spectral data were extracted for each IFU pixel within the vicinity of each emission band, and a polynomial continuum fit was performed, excluding the spectral region directly inside each molecular emission band (see Section 4). For CO2 and CO, a 3rd-order polynomial was used, whereas for H2O a 5th-order polynomial was used, to better fit the wing of the 3.0 m ice band . After continuum subtraction, the IFU spectra were integrated across the detected full emission width of each feature, then plotted as maps in Figure 2. The spatially averaged, continuum-subtracted emission band profile for each molecule is shown in the upper-right inset of their respective panels. The m scattered light image (integrated between 0.8–1.4 m) is also shown. The continuum brightness at 0.75 and 1.25 µm is 1013 and 1244 Jy inside a -diameter circular aperture, corresponding to cometary quantities (M. F. A’Hearn et al., 1984) of 392 and 492 cm, respectively (quoted without a correction for phase darkening).
@@
The gas and dust maps for 3I/ATLAS show a well-defined peak, offset East from the center of the IFU by (Figure 2). As shown by a later ephemeris reconstruction, about of the offset can be explained by a error on the predicted ephemeris, while the remaining offset remains under investigation; the astrometry available to-date shows no evidence for non-gravitational acceleration of 3I/ATLAS. The 1.2 m scattered light and CO2 emission maps reveal an extended coma of dust and gas that spans the full extent of the NIRSpec IFU. The H2O and CO emission is weaker, and therefore noisier, but nevertheless confirms the presence of a spatially extended molecular coma. While the scattered light shows a clear asymmetry along the Sun-comet axis — enhanced in the direction of the sky-projected comet-Sun and velocity vectors (see also D. Jewitt et al. 2025; C. O. Chandler et al. 2025) — the gas distributions (particularly CO2 and CO), appear relatively more symmetrical.
@@
To further investigate the coma structure, a enhanced version of the dust map (where is the sky-projected nucleocentric distance), is shown in Figure 3, and similarly for the gas maps in Appendix A (Figure 7). In these enhanced maps, the dominant coma spatial feature: dilution of the observed column densities due to quasi-spherical expansion, has been divided out. The -enhanced 1.2 m map reveals a strong, plume-like feature emanating from the pseudo-nucleus in the approximate direction of the Sun (slightly north of west), with an additional, weaker feature to the north-east. Since the dust in relatively faint comets such as 3I/ATLAS is optically thin at this wavelength, the shape of this feature is interpreted as resulting from an enhanced coma dust density in the sunward direction, potentially from the fragmentation of dust grains increasing the scattering cross section with distance from the nucleus (M. R. Combi, 1994; T. J. Jones et al., 2008). The gas maps on the other hand, show more subtle asymmetries. CO2 exhibits an azimuthal minimum towards the north, that we attribute to a combination of weak coma sub-structure and optical depth effects, since CO2 becomes optically thick close to the nucleus.
@@
Taking the ratio of continuum-subtracted fluxes within a -radius circular aperture centered from the brightest pixel on the sunward () and antisunward () sides, gives for the (1.2 m) dust, 1.03 for CO2, 1.31 for H2O and 1.00 for CO (with uncertainties of on all measurements). Our observations thus reveal a heterogeneous coma morphology consistent with different outgassing patterns for the different species. Such heterogeneity can be explained by various factors, such as the different molecular sublimation temperatures, nucleus release and coma acceleration mechanisms, as well as differing inertial/fluid-dynamical properties for the gas and dust. The degree of asymmetry of the observed gases is likely related to their sublimation temperatures (), with , congruent with (M. Womack et al., 2017). This implies that the sublimation of CO2 and CO is more fully activated than H2O. Full interpretation of the observed outgassing morphologies, including the surprisingly strong sunward dust enhancement, will require detailed physical modeling. Nevertheless, our data are consistent with the origin of this dust feature being influenced by enhanced gas sublimation (and therefore, outgassing) rates associated with the higher temperature on the dayside of the nucleus.
@@
The 3.0 µm H2O ice band depth is 78.30.2% at 2.9–3.1 µm, with respect to the continuum at 2.5 and 3.8 µm. The band remains strong across the field of view, varying by no more than 10%. Water ice also has broad near-infrared absorption features at 1.5, 2.0, 3.0, and 4.5 µm. The 4.5 µm band is difficult to measure in our data due to blending with CO2 and CO gas emission and likely thermal continuum, the 2.0 µm band appears very weak, and there is no discernible 1.5 µm band. Furthermore, the 2.0 µm spectral region in our combined IFU dataset is affected by several bad pixels. The average 2.0 µm band depth in the two dither positions with clean spectra in this region is (1.80.1)%. This was estimated by normalizing the reflectance spectrum within a 04 radius aperture with a linear fit between 1.75 and 2.22 µm, and measuring the mean value at 1.95–2.05 µm. The appearance of a strong 3-µm band, but weak or absent 1.5- and 2.0-µm bands implies the ice grains are micrometer size or smaller, while the shape of the band may be consistent with contributions from crystalline as well as amorphous ice (S. Protopapa et al., 2014).
@@ 4 Spectral Modeling
To derive gas production rates () and rotational temperatures (), the background-subtracted IFU data for CO2, CO, and H2O were subject to spectral modeling using optimal estimation routines as part of the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018). Initially, we constructed spectral models by taking the average spectrum for each gas inside a (1240 km) circular aperture centered on the nucleus. The relatively high signal-to-noise ratio (SNR) in this IFU region assisted in helping define the choice of continuum shape and the gases to be included in the model. Figure 4 shows the integrated spectra within this aperture, along with the best-fitting PSG model gas emission components. For the m region, we also tentatively identified an emission band at 4.85 m consistent with OCS, which was included in the fit. Additional details on the continuum fitting and spectral modeling procedure are given in Appendix B.
After modeling the nucleus-centered IFU extract, we proceeded to derive production rates and rotational temperatures as a function of , by extracting and modeling the data within successive -wide annular sectors, as shown in the inset diagram of Figure 5. The resulting production rates as a function of (referred to as “ curves”) are also shown in Figure 5
@@
The curves for CO2 and CO (Figure 5) level off towards larger nucleocentric distances, indicating that gas production for these species is confined within km of the nucleus. On the other hand, the H2O -curve shows no clear asymptote, and the error bars allow for the possibility of continued H2O production towards the edge of the NIRSpec IFU. Therefore, we cannot rule out a contribution to H2O gas in the outer coma from sublimating icy grains, which may be expected based on our detection of coma H2O ice. Our terminal value is significantly smaller than the value of s-1 measured at au by Z. Xing et al. (2025). This could be due to a rapid increase in H2O production between –2.9 au, with a possible contribution from icy-grain sublimation within the relatively large, 20,000 km diameter Swift aperture.
Although 13CO2 is securely detected in the nucleus-centered extract (Figure 4), optical depth effects and blending with the 12CO2 wing preclude the derivation of a reliable 13CO2 production rate in this region. We attempted to retrieve (13CO2) in the annular sector between – from the nucleus, but only an upper limit s-1 could be obtained, resulting in a () lower limit on the 12C/13C ratio of , which is formally consistent with the terrestrial value of 89 (T. B. Coplen et al., 2002). See Appendix B for further details.